# Beware of the Classical Benchmark Instances for the Traveling Salesman Problem with Time Windows

### v1

@@ 3 A Simple Solver for the TSPTW-M
The solver for the TSPTW-M builds upon the best-first search algorithm depicted in Algorithm 1. Given $t_{0},ub\in[T]$, Algorithm 1 determines whether an elementary route $R$ in the transport network $D$ is feasible and arrives earlier than $ub$ when departing at time $t_{0}$, i.e., whether $\delta(R,t_{0})\leq ub$. If $t_{0}=0$, then the makespan of $R$ is less than $ub$.
@@
In essence, Algorithm 1 is an informed search procedure that traverses a tree of partial routes $\mathcal{T}$ in a backward direction. The root of $\mathcal{T}$ is the partial route $\langle n+1\rangle$, which only visits the end depot, while the children of a partial route $R$ are all the elementary partial routes $\langle w\rangle+R$ obtained by prepending a vertex $w\not\in V(R)$ to $R$. Here, $V(R)$ denotes the set of vertices set already visited by $R$.
Within Algorithm 1, $\mathcal{R}$ maintains the unprocessed nodes of $\mathcal{T}$, which have already had their parent processed (Steps 3 and 14). Since $\mathcal{T}$ contains $\Theta(n!)$ nodes, two techniques are employed in Step 14 to prevent the exploration of the entire tree. First, a set of unreachable vertices $U(w,\delta^{-1}_{\rm m}(R))$ is computed for every partial route $R$ that starts at vertex $w$, as described in Section 3.1. By construction, if $U(w,\delta^{-1}_{\rm m}(R))\not\subseteq V(R)$, then no elementary route $R^{\prime}$ with suffix $R$ and $\delta(R^{\prime})<ub$ exists, and therefore the subtree of $\mathcal{T}$ rooted at $R$ can be discarded. Similarly, if $\delta_{\rm m}(R)\geq ub$, then $\delta_{\rm m}(R^{\prime})\geq ub$ for every route $R^{\prime}$ with suffix $R$, and so this subtree can also be pruned. These pruning strategies ensure that any partial route $R$ obtained in Step 8 is feasible and has $\delta_{\rm m}(R)<ub$.
Following FontaineDibangoyeSolnon2023, each iteration of the main loop (Steps 414) processes one route $R$ of length $k$, if one exists, for every $k\in[\![n+1]\!]$ (Steps 514). All extensions of $R$ in Step 14 will have length $k+1$, and thus can be processed within the same iteration of the main loop. The goal is to compute the output route as early as possible. To guide the search, a heuristic is applied to decide which partial route $R$ of length $k$ to process next (Step 7).
Note that $ub-\delta_{\rm m}(R)$ represents an upper bound on the time that can still be saved by any route with suffix $R$. In the extreme case where $\delta_{\rm m}(R)\geq ub$, no improvement is possible, and $R$ is discarded (Step 14). Maximizing $ub-\delta_{\rm m}(R)$ maximizes the opportunity for the vehicle to wait at previous customers, if necessary. In event of ties, Step 7 minimizes $\delta^{-1}(R,\delta_{\rm m}(R))$ to maximize the earliest departure time of $R$ that yields no waiting due to time windows. The goal is to allow the vehicle to wait at the previous (yet unknown) customers without significantly affecting the earliest arrival time.
@@
5:  for all $k\in[\![n+1]\!]$:
6:   if $\mathcal{R}$ has no route of length $k$: continue
7:   Remove from $\mathcal{R}$ a partial route $R$ of length $k$ minimizing $\langle\delta_{\rm m}(R),\delta^{-1}(R,\delta_{\rm m}(R))\rangle$
@@
13:   else: let $\operatorname{best}(V(R),v)=\langle\delta^{-1}_{\rm m}(R),\delta(R,\delta^{-1}_{\rm m}(R))\rangle$
14:   for all $w\in[\![n]\!]-V(R)$: insert $R^{\prime}=\langle w\rangle+R$ into $\mathcal{R}$ if $\delta_{\rm m}(R^{\prime})<ub$ and $U(w,\delta^{-1}_{\rm m}(R^{\prime}))\subseteq V(R^{\prime})$
15:output “no route $R$ exists with $\delta_{\rm m}(R)<ub$
@@
3:let $R^{*}=\emptyset$ and $T^{\prime}=T+1$.
4:Compute the unreachable function for Step 14 as in Section 3.1
@@
Implementation details.
Although our implementation of Algorithm 1 is not particularly efficient, we have taken measures to improve its running time. Each partial route $R$ is implemented with a label $\ell(R)$ that uses $O(1)$ bits. The label $\ell(R)$ is a tuple containing the following information:
@@
  • The other limiting times $\delta(R,\delta^{-1}_{\rm m}(R,ub))$ and $\delta^{-1}(R,\delta_{\rm m}(R))$, and
@@
By traversing the branch of $\mathcal{T}$ from node $R$ using the parents pointers, we can obtain all the vertices in $R$ in $O(n)$ time. Each time $R$ is extended into a a partial route $R^{\prime}$ in Step 14, the times $\delta(R^{\prime},\delta^{-1}_{\rm m}(R^{\prime}))$ and $\delta^{-1}(R^{\prime},\delta_{\rm m}(R^{\prime}))$ are computed in $O(n)$ time. Additionally, $U(w,\delta^{-1}_{\rm m}(R^{\prime}))$ is computed in $O(n)$ time, as discussed in Section 3.1.
@@ 3.1 The Unreachable Function
As illustrated in Algorithm 2 (Step 4), we apply a preprocessing step to compute an unreachable function $U\colon[\![n+1]\!]\times[T]\to 2^{[\![n+1]\!]}$, which is later used by Algorithm 1 to prune some branches of the tree $\mathcal{T}$ of partial routes. For each vertex $v\in[\![n+1]\!]$ and time $t\in[T]$, the function $U(v,t)$ returns a set of vertices that cannot appear before $v$ in any elementary and feasible route $R$ that visits $v$ at a time $t^{\prime}\leq t$. By construction, $U(v,t)\subseteq U(v,t^{\prime})$, so $U$ is implemented by storing $U(v,t)$ only for those times $t\in[T]$ where the function changes. As a result, each $U(v,\bullet)$ has $O(n)$ values, which allows for queries to be answered in $O(n)$ time.
@@
[…]
@@
$\operatorname{EAT}(z,w)>\operatorname{LDT}(w,v),\quad\operatorname{EAT}(w,v)>\operatorname{LDT}(v,z),\quad\text{and}\quad\operatorname{EAT}(w,z)>\operatorname{LDT}(z,v).$
@@ 3.2 Solving the integer TSPTW-D
Recall that the TSPTW-D consist of finding an elementary route $R$ with minimum duration $\Delta(R)=\min\{\delta(R,t)-t\mid t\in[T]\}$ for a given transport network $D$. Clearly, if the travel times and all the time windows in $D$ are integers, then a solution to the TSPTW-D can be obtained by running a solver for the TSPTW-M $O(T)$ times. This is the approach Algorithm 3 uses to solve the TSPTW-D.
Specifically, Algorithm 3 maintains a sliding window [[p,q]][\![p,q]\!], a route $R$ that minimizes $\delta(R,p-1)=q$, and a route $R^{*}$ with a time $t^{*}$ that minimizes $\delta(R^{*},t^{*})-t^{*}$ for $t^{*}\in[\![p-1]\!]$. Steps 35 initialize these variables. Step 5 also computes $\operatorname{last}_{p}$, which is the latest of the departure times of all the elementary routes in $D$. Thus, when $p>\operatorname{last}_{p}$, $R^{*}$ is a route with minimum duration.
Throughout the main loop (Steps 611) the route $R$ is updated into a new route that minimizes $\delta(R,p)\geq q$. Note that $\delta(R,p)=q$ when Step 9 is skipped. Then, $q$, $R^{*}$, and $t^{*}$ are updated accordingly in Steps 9, 11, and 12.
@@
2:a route $R^{*}$ and a time $t^{*}$ minimizing $\Delta(R^{*})<T$ or a message that $D$ is infeasible.
3:let $R$ be a route minimizing $\delta(R)$; halt with outputDD is infeasible” if $\delta(R)=\infty$
4:Heuristically find a route $R^{*}$ and a time $t^{*}$ with $\delta(R^{*},t^{*})=\delta(R)$ maximizing $t^{*}$
5:let $p=t^{*}+1$, $q=\delta(R^{*})$, and $\operatorname{last}_{p}=\max\{\delta^{-1}(R,T)\mid R\text{ is an elementary route}\}$
6:while $p\leq\operatorname{last}_{p}$:
7:  Heuristically update $R$ to minimize $\delta(R,p)$
8:  if $\delta(R,p)>q$:
9:   let $R$ be a route minimizing $\delta(R,p)$ and let $q=\delta(R,p)$
10:  Heuristically update $R$ and $p$ to maximize $p$ with $\delta(R,p)=q$
11:  if $\delta(R^{*},t^{*})-t^{*}>\delta(R,p)-p$: let $R^{*}=R$ and $t^{*}=p$
12:  let $p=p+1$
13:output $R^{*}$ and $t^{*}$
Regarding the implementation of Algorithm 3, we run Algorithm 2 with $t_{0}=0$ and $ub=T+1$ for Step 3. Similarly, we invoke Algorithm 2 with $t_{0}=p$ and $ub=\min\{\operatorname{last}_{q}+\Delta(R^{*}),T+1\}$ for Step 9. Let $D^{-1}$ be the reverse transport network with vertex set [[n+1]][\![n+1]\!], where the travel time of an arc $v\to w$ is $\tau(w,v)$, and the time window of a vertex $v$ is [Tb(v),Ta(v)][T-b(v),T-a(v)]. In $D^{-1}$, the start depot is $n+1$ and the end depot is 0. To initialize $\operatorname{last}_{p}$ in Step 5, we run Algorithm 2 on $D^{-1}$ with $t_{0}=0$ and $ub=T-t^{*}$. If a route $R^{-1}$ in $D^{-1}$ is found, then $\operatorname{last}_{p}=T-\delta(R^{-1},0)$; otherwise, $\operatorname{last}_{p}=t^{*}$. Finally, we use a local search heuristic for Steps 7 and 10, which applies the swap, $2$-opt, and shift operators with a first-improvement strategy. Our implementation is not particularly efficient, as it spends $O(n^{2})$ time on each operator.
@@ 4.1 Results for the TSPTW-M
For the TSPTW-M we set a time limit of $3$ minutes per instance. Although this time limit is quite restrictive for an exact method, it is sufficient for our purposes and helps prevent unnecessary computation time before encountering a memory limit exception. Furthermore, this time limit is representative of a typical use case, where an enumerative algorithm is applied for a short duration after computing a tight lower bound (see e.g. BaldacciMingozziRoberti2012; TilkIrnich2017; Lera-RomeroMirandaBrontSoulignac2022).
The results for the classical benchmarks and DaS are summarized in Table 1. We compare Algorithm 2 against the methods by Lera-RomeroMirandaBrontSoulignac2022 (Ler22), RudichCappartRousseau2023 (Rud23), and FontaineDibangoyeSolnon2023 (Fon23), which represent the state of the art for the TSPTW-M within exact methods. Ler22 and Fon23 were originally proposed for time-dependent versions of the TSPTW, and thus Lera-RomeroMirandaBrontSoulignac2022 and FontaineDibangoyeSolnon2023 did not report the results for the TSPTW-M. Similarly, RudichCappartRousseau2023 did not test Rud23 on the DaS benchmark. These missing results reported in Table 1 are taken from Fontaine2023. We note that the results in Fontaine2023 differ from those reported by RudichCappartRousseau2023 because Fontaine2023 executed all the algorithms on the same machine and with the same configuration. Our computer has a slightly higher performance for single thread execution according to www.cpubenchmark.net, and, unlike Fontaine2023, we did not disable turbo mode (we kept the ondemand governor within Ubuntu). Additionally, the time limit considered by Fontaine2023 was one hour per instance. These discrepancies do not affect our main conclusions, and we thus disregard them.
In Table 1, column # counts the number of instances in each benchmark set, columns $s$ indicate the number of solved instances, and columns $t_{s}$ report the average time for the solved instances, rounded to the nearest second. For Algorithm 2, we also include the maximum running time among the solved instances in column $m_{s}$. Additionally, we divide the results for Asc into Asc-s and Asc-l to separately report the results for instances with fewer than $50$ customers and those with $50$ or more customers, respectively. Table 1 reports the results for Rud23 with the unseeded version outside parentheses and the seeded version inside parentheses. The difference between the two is that the latter is initialized with a known solution. In this regard, Algorithm 2 should be considered unseeded. Similarly, Table 1 reports the results for Fon23 for two versions of the heuristic guiding the informed search. The results for FEA are reported outside parentheses, and those for MSA are reported inside parentheses.
Ler22 Rud23 Fon23 Algorithm 2
@@
Table 1: Results of Algorithm 2 on the classical benchmarks. Entries marked with * were reported by Fontaine2023.
@@
A quick look at Table 1 is enough to conclude that Algorithm 2 outperforms the state-of-the-art methods on the larger instances, solving the four instances that remained unsolved. Although this is strictly true, knowing the inner workings of Algorithm 2, we would never prefer it as a generic exact solver for the TSPTW-M. A deeper look at Table 1 raises an important concern: Algorithm 2 is unable to solve seven instances with 45 or fewer customers (and our unreported experiments show that memory is exhausted when the time limit is high enough). Furthermore, Algorithm 2 performs better on the larger instances in Asc than on the smaller ones. As strange as it may seem, RifkiSolnon2025 already noted that the ratio between the size of the time windows and the horizon could be more important than the number of customers when evaluating the efficiency of dynamic programming solvers for the TSPTW-M. Finally, we observe that Algorithm 2 is particularly fast on Lan, Dum, Gen, Ohl, and DaS, all of which follow the same creation procedure commonly applied in many machine learning methods to create “hard” training datasets (Section 2.1).
@@
What we believe is happening is that Algorithm 2 takes advantage of some inherent bias in the benchmark instances resulting from the tight time windows, which yield a small search space of feasible solutions that have a well defined structure. Fon23 is another informed search that seems to have an advantage because of the structure of the instances, as depicted by the Ohl benchmark: the $20$ instances it solves are completed in just a few seconds, while Ler22 is unable to solve any of them within an hour! The difference is that Ler22 was designed to solve as many as possible of the small but challenging instances. As a result, it spends a large amount of time computing penalties on a relaxed problem before running its enumeration algorithm, and it uses a pure best-first search heuristic to guide its informed search, avoiding the extension of dominated labels. Finally, Rud23 is a generic algorithm based on multivalued decision diagrams that iteratively traverses enumeration trees of different widths to obtain lower and upper bounds to “peel” the tree. Being able to find feasible solutions quickly improves the peeling process.
@@
To present the contrasting results, we report the performance of the algorithms for the Rif benchmark in Table 2. (The results for Ler22, Rud23, and Fon23 were taken from Fontaine2023). Again, Algorithm 2 is among the fastest for tight time windows ($\beta\in\{50,100\}$), but it is clearly the worst for the looser time windows ($\beta\in\{0,25\}$), even when compared against the generic method Rud23. (We remark that increasing the time limit does not help in these cases, as the memory limit is always reached.) The clear winner in robustness is Ler22, which was specifically designed for harder instances in terms of time windows.
$\beta$ $n$ Ler22 Rud23 Fon23 Algorithm 2
@@
In our opinion, the main conclusion from the experiments in this section is that the classical benchmarks are no longer adequate for evaluating the TSPTW-M, as they can be “exploited” by a simple algorithm. Consequently, any method that incorporates a variant of Algorithm 2 in its preprocessing step would likely produce results that appear outstanding to an external reviewer.

### v2

@@ 3 A Simple Solver for the TSPTW-M
The solver for the TSPTW-M builds upon the best-first search algorithm depicted in Algorithm 1. Given $t_{0},ub\in[T]$, Algorithm 1 determines whether a route $R$ in the transport network $D$ is feasible and arrives earlier than $ub$ when departing at time $t_{0}$, i.e., whether $\delta(R,t_{0})\leq ub$. If $t_{0}=0$, then the makespan of $R$ is less than $ub$.
@@
In essence, Algorithm 1 is an informed search procedure that traverses a tree of partial routes $\mathcal{T}$ in a backward direction. The root of $\mathcal{T}$ is the partial route $\langle n+1\rangle$, which only visits the end depot, while the children of a partial route $R$ are all the partial routes $\langle w\rangle+R$ obtained by prepending a vertex $w\not\in V(R)$ to $R$. Here, $V(R)$ denotes the set of vertices set already visited by $R$.
Within Algorithm 1, $\mathcal{R}$ maintains the unprocessed nodes of $\mathcal{T}$, which have already had their parent processed (Steps 3 and 15). Since $\mathcal{T}$ contains $\Theta(n!)$ nodes, two techniques are employed in Step 16 to prevent the exploration of the entire tree. First, a set of unreachable vertices $U(w,\delta^{-1}_{\rm m}(R))$ is computed for every partial route $R$ that starts at vertex $w$, as described in Section 3.1. By construction, if $U(w,\delta^{-1}_{\rm m}(R))\not\subseteq V(R)$, then no route $R^{\prime}$ with suffix $R$ and $\delta(R^{\prime})<ub$ exists, and therefore the subtree of $\mathcal{T}$ rooted at $R$ can be discarded. Similarly, if $\delta_{\rm m}(R)\geq ub$, then $\delta_{\rm m}(R^{\prime})\geq ub$ for every route $R^{\prime}$ with suffix $R$, and so this subtree can also be pruned. These pruning strategies ensure that any partial route $R$ obtained in Step 8 is feasible and has $\delta_{\rm m}(R)<ub$.
Following Fontaine et al. (2023), each iteration of the main loop (Steps 416) processes one route $R$ of length $k$, if one exists, for every $k\in[\![n+1]\!]$ (Steps 516). All extensions of $R$ in Step 15 will have length $k+1$, and thus can be processed within the same iteration of the main loop. The goal is to compute the output route as early as possible. To guide the search, a heuristic is applied to decide which partial route $R$ of length $k$ to process next (Step 7).
Note that $ub-\delta_{\rm m}(R)$ represents an upper bound on the time that can still be saved by any route with suffix $R$. In the extreme case where $\delta_{\rm m}(R)\geq ub$, no improvement is possible, and $R$ is discarded (Step 16). Maximizing $ub-\delta_{\rm m}(R)$ maximizes the opportunity for the vehicle to wait at previous customers, if necessary. In event of ties, Step 7 minimizes $\delta^{-1}(R,\delta_{\rm m}(R))$ to maximize the earliest departure time of $R$ that yields no waiting due to time windows. The goal is to allow the vehicle to wait at the previous (yet unknown) customers without significantly affecting the earliest arrival time.
@@
5:  for all $k=0,\ldots,n+1$:
6:   if $\mathcal{R}$ has no partial route of length $k$: continue
7:   Remove $R\in\mathcal{R}$ of length $k$ and minimum $\langle\delta_{\rm m}(R),\delta^{-1}(R,\delta_{\rm m}(R))\rangle$
@@
13:   let $\operatorname{best}(V(R),v)=\langle\delta^{-1}_{\rm m}(R),\delta(R,\delta^{-1}_{\rm m}(R))\rangle$
14:   for all $w\in[\![n]\!]-V(R)$:
15:     let $R^{\prime}=\langle w\rangle+R$
16:     if $\delta_{\rm m}(R^{\prime})<ub$ and $U(w,\delta^{-1}_{\rm m}(R^{\prime}))\subseteq V(R^{\prime})$: insert $R^{\prime}$ into $\mathcal{R}$
17:output “no route $R$ exists with $\delta_{\rm m}(R)<ub$
@@
3:compute the unreachable function for Step 16 as in Section 3.1
4:let $R^{*}=\emptyset$ and $T^{\prime}=T+1$.
@@
Implementation details
Although our implementation of Algorithm 1 is not particularly efficient, we have taken measures to improve its running time. Each partial route $R$ is implemented with a label $\ell(R)$ that uses $n+O(1)$ bits. The label $\ell(R)$ is a tuple containing the following information:
@@
  • The times $\delta(R,\delta^{-1}_{\rm m}(R))$ and $\delta^{-1}(R,\delta_{\rm m}(R))$, and
@@
By traversing the branch of $\mathcal{T}$ from node $R$ using the parents pointers, we can obtain all the vertices in $R$ in $O(n)$ time. Each time $R$ is extended into a partial route $R^{\prime}$ in Step 15, the times $\delta(R^{\prime},\delta^{-1}_{\rm m}(R^{\prime}))$ and $\delta^{-1}(R^{\prime},\delta_{\rm m}(R^{\prime}))$ are computed in $O(n)$ time. Additionally, $U(w,\delta^{-1}_{\rm m}(R^{\prime}))$ is computed in $O(n)$ time, as discussed in Section 3.1.
@@ 3.1 The Unreachable Function
As illustrated in Algorithm 2 (Step 3), we apply a preprocessing step to compute an unreachable function $U\colon[\![n+1]\!]\times[T]\to 2^{[\![n+1]\!]}$, which is later used by Algorithm 1 to prune some branches of the tree $\mathcal{T}$ of partial routes. For each vertex $v\in[\![n+1]\!]$ and time $t\in[T]$, the function $U(v,t)$ returns a set of vertices that cannot appear before $v$ in any feasible route $R$ that visits $v$ at a time $t^{\prime}\leq t$. By construction, $U(v,t)\subseteq U(v,t^{\prime})$, so $U$ is implemented by storing $U(v,t)$ only for those times $t\in[T]$ where the function changes. As a result, each $U(v,\bullet)$ has $O(n)$ values, which allows for queries to be answered in $O(n)$ time.
@@
$\operatorname{LDT}(w,v)$ $=\max\{\delta^{-1}(R,b(v))\mid R\text{ is a feasible from }v\text{ to }w\}\cup\{-\infty\}$
@@
$\operatorname{EAT}(z,w)$ $>\operatorname{LDT}(w,v),$
$\operatorname{EAT}(w,v)$ $>\operatorname{LDT}(v,z),\text{and}$
$\operatorname{EAT}(w,z)$ $>\operatorname{LDT}(z,v).$
@@ 3.2 Solving the integer TSPTW-D
Recall that the TSPTW-D consist of finding a route $R$ with minimum duration $\Delta(R)=\min\{\delta(R,t)-t\mid t\in[T]\}$ for a given transport network $D$. Clearly, if the travel times and all the time windows in $D$ are integers, then a solution to the TSPTW-D can be obtained by running a solver for the TSPTW-M $O(T)$ times. This is the approach Algorithm 3 uses to solve the TSPTW-D.
Specifically, Algorithm 3 maintains a sliding window [[p,q]][\![p,q]\!], a route $R$ that minimizes $\delta(R,p-1)=q$, and a route $R^{*}$ with a time $t^{*}$ that minimizes $\delta(R^{*},t^{*})-t^{*}$ for $t^{*}\in[\![p-1]\!]$. Steps 36 initialize these variables. Step 6 also computes $\operatorname{last}_{p}$, which is the latest of the departure times of all the routes in $D$. Thus, when $p>\operatorname{last}_{p}$, $R^{*}$ is a route with minimum duration.
Throughout the main loop (Steps 712) the route $R$ is updated into a new route that minimizes $\delta(R,p)\geq q$. Note that $\delta(R,p)=q$ when Step 10 is skipped. Then, $q$, $R^{*}$, and $t^{*}$ are updated accordingly in Steps 10, 12, and 13.
@@
2:a route $R^{*}$ and a time $t^{*}$ minimizing $\Delta(R^{*})=\delta(R,t^{*})-t^{*}$ or a message that $D$ is infeasible.
3:let $R$ be a route minimizing $\delta(R)$
4:halt with outputDD is infeasible” if $\delta(R)=\infty$
5:heuristically find a route $R^{*}$ with $\delta(R^{*},t^{*})=\delta(R)$ that maximizes $t^{*}\in[T]$
6:let $p=t^{*}+1$, $q=\delta(R^{*})$, and $\operatorname{last}_{p}=\max\{\delta^{-1}(R,T)\mid R\text{ is a route in }D\}$
7:while $p\leq\operatorname{last}_{p}$:
8:  heuristically update $R$ to minimize $\delta(R,p)$
9:  if $\delta(R,p)>q$:
10:   let $R$ be a route that minimizes $\delta(R,p)$ and let $q=\delta(R,p)$
11:  heuristically update $R$ and $p$ to maximize $p$ with $\delta(R,p)=q$
12:  if $\delta(R^{*},t^{*})-t^{*}>\delta(R,p)-p$: let $R^{*}=R$ and $t^{*}=p$
13:  let $p=p+1$
14:output $R^{*}$ and $t^{*}$
Regarding the implementation of Algorithm 3, we run Algorithm 2 with $t_{0}=0$ and $ub=T+1$ for Step 3. Similarly, we invoke Algorithm 2 with $t_{0}=p$ and $ub=\min\{\operatorname{last}_{q}+\Delta(R^{*}),T+1\}$ for Step 10. Let $D^{-1}$ be the reverse transport network with vertex set [[n+1]][\![n+1]\!], where the travel time of an arc $v\to w$ is $\tau(w,v)$, and the time window of a vertex $v$ is [Tb(v),Ta(v)][T-b(v),T-a(v)]. In $D^{-1}$, the start depot is $n+1$ and the end depot is 0. To initialize $\operatorname{last}_{p}$ in Step 6, we run Algorithm 2 on $D^{-1}$ with $t_{0}=0$ and $ub=T-t^{*}$. If a route $R^{-1}$ in $D^{-1}$ is found, then $\operatorname{last}_{p}=T-\delta(R^{-1},0)$; otherwise, $\operatorname{last}_{p}=t^{*}$. Finally, we use a local search heuristic for Steps 8 and 11, which applies the swap, $2$-opt, and shift operators with a first-improvement strategy. Our implementation is not particularly efficient, as it spends $O(n^{2})$ time on each operator.
@@ 4.1 Results for the TSPTW-M
For the TSPTW-M we set a time limit of $3$ minutes per instance. Although this time limit is quite restrictive for an exact method, it is sufficient for our purposes and helps prevent unnecessary computation time before encountering a memory limit exception. Furthermore, this time limit is representative of a typical use case, where an enumerative algorithm is applied for a short duration after computing a tight lower bound (see e.g. Baldacci et al., 2012; Tilk and Irnich, 2017; Lera-Romero et al., 2022).
The results for the classical benchmarks and DaS are summarized in Table 1. We compare Algorithm 2 against the methods by Lera-Romero et al. (2022) (Ler22), Rudich et al. (2023) (Rud23), and Fontaine et al. (2023) (Fon24), which represent the state of the art for the TSPTW-M within exact methods. Ler22 and Fon24 were originally proposed for time-dependent versions of the TSPTW, and thus Lera-Romero et al. (2022) and Fontaine et al. (2023) did not report the results for the TSPTW-M. Similarly, Rudich et al. (2023) did not test Rud23 on the DaS benchmark. These missing results reported in Table 1 are taken from Fontaine (2024). We note that the results in Fontaine (2024) differ from those reported by Rudich et al. (2023) because Fontaine (2024) executed all the algorithms on the same machine and with the same configuration. Our computer has a slightly higher performance for single thread execution according to www.cpubenchmark.net, and, unlike Fontaine (2024), we did not disable turbo mode (we kept the ondemand governor within Ubuntu). Additionally, the time limit considered by Fontaine (2024) was one hour per instance. These discrepancies do not affect our main conclusions, and we thus disregard them.
In Table 1, column # counts the number of instances in each benchmark set, columns $s$ indicate the number of solved instances, and columns $t_{s}$ report the average time for the solved instances, rounded to the nearest second. For Algorithm 2, we also include the maximum running time among the solved instances in column $m_{s}$. Additionally, we divide the results for Asc into Asc-s and Asc-l to separately report the results for instances with fewer than $50$ customers and those with $50$ or more customers, respectively. Table 1 reports the results for Rud23 with the unseeded version outside parentheses and the seeded version inside parentheses. The difference between the two is that the latter is initialized with a known solution. In this regard, Algorithm 2 should be considered unseeded. Similarly, Table 1 reports the results for Fon24 for two versions of the heuristic bound guiding the informed search: FEA (feasiblity) bound and MSA (minimum spanning arborescence) bound. The results for FEA are reported outside parentheses, and those for MSA are reported inside parentheses.
Ler22 Rud23 Fon24 Algorithm 2
@@
Table 1: Results of Algorithm 2 on the classical benchmarks. Entries marked with * were reported by Fontaine (2024).
@@
A quick look at Table 1 is enough to conclude that Algorithm 2 outperforms the state-of-the-art methods on the larger instances, solving the four instances that remained unsolved. Although this is strictly true, knowing the inner workings of Algorithm 2, we would never prefer it as a generic exact solver for the TSPTW-M. A deeper look at Table 1 raises an important concern: Algorithm 2 is unable to solve seven instances with 45 or fewer customers (and our unreported experiments show that memory is exhausted when the time limit is high enough). Furthermore, Algorithm 2 performs better on the larger instances in Asc than on the smaller ones. As strange as it may seem, Rifki and Solnon (2025) already noted that the ratio between the size of the time windows and the horizon could be more important than the number of customers when evaluating the efficiency of dynamic programming solvers for the TSPTW-M. Finally, we observe that Algorithm 2 is particularly fast on Lan, Dum, Gen, Ohl, and DaS, all of which follow the same creation procedure commonly applied in many machine learning methods to create “hard” training datasets (Section 2.1).
@@
What we believe is happening is that Algorithm 2 takes advantage of some inherent bias in the benchmark instances resulting from the tight time windows, which yield a small search space of feasible solutions that have a well defined structure. Fon24 is another informed search that seems to have an advantage because of the structure of the instances, as depicted by the Ohl benchmark: the $20$ instances it solves are completed in just a few seconds, while Ler22 is unable to solve any of them within an hour! The difference is that Ler22 was designed to solve as many as possible of the small but challenging instances. As a result, it spends a large amount of time computing penalties on a relaxed problem before running its enumeration algorithm, and it uses a pure best-first search heuristic to guide its informed search, avoiding the extension of dominated labels. Finally, Rud23 is a generic algorithm based on multivalued decision diagrams that iteratively traverses enumeration trees of different widths to obtain lower and upper bounds to “peel” the tree. Being able to find feasible solutions quickly improves the peeling process.
@@
To present the contrasting results, we report the performance of the algorithms for the Rif benchmark in Table 2. (The results for Ler22, Rud23, and Fon24 were taken from Fontaine, 2024). Again, Algorithm 2 is among the fastest for tight time windows ($\beta\in\{50,100\}$), but it is clearly the worst for the looser time windows ($\beta\in\{0,25\}$), even when compared against the generic method Rud23. (We remark that increasing the time limit does not help in these cases, as the memory limit is always reached.) The clear winner in robustness is Ler22, which was specifically designed for harder instances in terms of time windows.
$\beta$ $n$ Ler22 Rud23 Fon24 Algorithm 2
@@
In our opinion, the main conclusion from the experiments in this section is that the classical benchmarks are no longer adequate when used as the sole tool for evaluating the TSPTW-M. If not complemented by other benchmarks, any method that incorporates a variant of Algorithm 2 in its preprocessing step would likely produce results that may appear outstanding to an external reviewer.